# The Design of an LLM-powered Unstructured Analytics System

### v1

@@ 4. Partitioner
The Aryn Partitioner can also identify and process images in documents using multi-modal LLMs. This allows for image metadata and textual summarization of image contents to be extracted and queried. Likewise, many enterprise documents contain images of printed or handwritten text, requiring an OCR stap for extracting the relevant text. This requires an accurate segmentation model to identify portions of the image that contain text, which can then be processed using OCR models such as EasyOCR [^5] https://github.com/JaidedAI/EasyOCR.
@@ 6. Luna
Luna enables users to ask complex questions on data without having to understand the underlying primitives and execution model. This approach works well for data exploration and simple questions, but as the nature of questions becomes more complex and the semantic nature of data comes into play, a fully automated system falls short. Lack of visibility into how an answer was computed results in a lack of trust in a query system. The inability to correct or refine a query causes significant difficulty in trying to get a good answer. Luna solves this by exposing a logical query execution plan, data lineage, and execution history for all queries. This allows a user to understand the exact operations that were performed to answer a query, how the dataset was transformed during each operation, and modify any part of the plan to better align with their query’s intention. With a combination of automated query undestanding, and efficient execution with full user control, we believe Luna is able to solve many real-world scenarios that data engineers and scientists face daily.
@@ 6.1. Luna architecture
Data schema. Luna operates on data ingested using Sycamore, benefitting from structured information extracted from unstructured data. Luna uses this schema during the query planning phase to determine the appropriate set of operators to use to answer a query. The schema can evolve over time, based on new semantic relationships discovered in the data.
Natural language inferface. Luna supports a natural language query interface, making it easy for users to get started without learning a new programming interface. Luna also expresses the query plans it produces as natural language text, making it easy for users to understand how Luna arrived at an answer.
@@
Plan optimization. Query operators vary significantly in latency, computational load, and monetary cost. The plan optimizer makes tradeoffs based on cost vs efficiency to handle different scenarios. It is able to combine and batch operations when possible, and make decisions about what technique (string matching vs semantic matching), and tool (e.g., GPT-4 versus Llama 7B) to use.
@@ 7. Related Work
There is less work on building end-to-end systems that encompass the entire spectrum of tasks from ETL to querying, but several efforts have started over the last year including ZenDB (Lin et al., 2024), LOTUS (Patel et al., 2024), EVAPORATE (Arora et al., 2023), CHORUS (Kayali et al., 2024), and Palimpzest (Liu et al., 2024a). In contrast to these works, we do not believe it is possible to fully automate the entire pipeline, and have designed Aryn to faciliate a human-in-the-loop paradigm.
The Aryn Partitioner is based on a long line of work in document segmentation. Current approaches commonly use object detection models such as DETR (Carion et al., 2020). The Aryn Partitioner follows this approach and leverages Deformable DETR (Zhu et al., 2020). An alternate line of work has led to multi-modal models such as Donut (Kim et al., 2021) and LayoutLMV3 (Huang et al., 2022) that seek to directly solve document understanding tasks like visual question answering (VQA) without the need for explicit segmentation. Sycamore can eventually incorporate these models, but we continue to find segementation valuable as we can index the segments to reduce work at query time.
@@ 8. Conclusions and Future Work
We are building Aryn to unify unstructured and structured analytics under a single abstraction, by leveraging the immense potential of LLMs to process multi-modal datasets. At the same time, given the limitations of current models, we are building Aryn to be a human-in-the-loop system; as the LLMs improve, the need for human interventions will diminish, but it is unlikely to completely vanish. Our experience across a variety of application domains has validated our overall design as well as Aryn’s individual components. Of course, many challeges still remain. We need more robust data and querying abstractions to support analytics across structured and unstructured repositories. Knowledge Graphs are emerging as a strong contender for this purpose, but it is not yet clear how hallucination-free knowledge graphs can be built in a cost-effective manner with minimum human intervention. We also need to build more scalable systems for processing large multi-modal repositories, for inferring connections between them, and for leveraging existing structured repositories, like data warehouses.

### v2

@@ 4. Partitioner
The Aryn Partitioner can also identify and process images in documents using multi-modal LLMs. This allows for image metadata and textual summarization of image contents to be extracted and queried. Likewise, many enterprise documents contain images of printed or handwritten text, requiring an OCR step for extracting the relevant text. This requires an accurate segmentation model to identify portions of the image that contain text, which can then be processed using OCR models such as EasyOCR [^5] https://github.com/JaidedAI/EasyOCR.
@@ 6. Luna
Luna enables users to ask complex questions on data without having to understand the underlying primitives and execution model. This approach works well for data exploration and simple questions, but as the nature of questions becomes more complex and the semantic nature of data comes into play, a fully automated system falls short. Lack of visibility into how an answer was computed results in a lack of trust in a query system. The inability to correct or refine a query causes significant difficulty in trying to get a good answer. Luna solves this by exposing a logical query execution plan, data lineage, and execution history for all queries. This allows a user to understand the exact operations that were performed to answer a query, how the dataset was transformed during each operation, and modify any part of the plan to better align with their query’s intention. With a combination of automated query understanding, and efficient execution with full user control, we believe Luna is able to solve many real-world scenarios that data engineers and scientists face daily.
@@ 6.1. Luna architecture
Data schema. Luna operates on data ingested using Sycamore, benefiting from structured information extracted from unstructured data. Luna uses this schema during the query planning phase to determine the appropriate set of operators to use to answer a query. The schema can evolve over time, based on new semantic relationships discovered in the data.
Natural language interface. Luna supports a natural language query interface, making it easy for users to get started without learning a new programming interface. Luna also expresses the query plans it produces as natural language text, making it easy for users to understand how Luna arrived at an answer.
@@
Plan optimization. Query operators vary significantly in latency, computational load, and monetary cost. The plan optimizer makes trade-offs based on cost vs efficiency to handle different scenarios. It is able to combine and batch operations when possible, and make decisions about what technique (string matching vs semantic matching), and tool (e.g., GPT-4 versus Llama 7B) to use.
@@ 7. Related Work
There is less work on building end-to-end systems that encompass the entire spectrum of tasks from ETL to querying, but several efforts have started over the last year including ZenDB (Lin et al., 2024), LOTUS (Patel et al., 2024), EVAPORATE (Arora et al., 2023), CHORUS (Kayali et al., 2024), and Palimpzest (Liu et al., 2024a). In contrast to these works, we do not believe it is possible to fully automate the entire pipeline, and have designed Aryn to facilitate a human-in-the-loop paradigm.
The Aryn Partitioner is based on a long line of work in document segmentation. Current approaches commonly use object detection models such as DETR (Carion et al., 2020). The Aryn Partitioner follows this approach and leverages Deformable DETR (Zhu et al., 2020). An alternate line of work has led to multi-modal models such as Donut (Kim et al., 2021) and LayoutLMV3 (Huang et al., 2022) that seek to directly solve document understanding tasks like visual question answering (VQA) without the need for explicit segmentation. Sycamore can eventually incorporate these models, but we continue to find segmentation valuable as we can index the segments to reduce work at query time.
@@ 8. Conclusions and Future Work
We are building Aryn to unify unstructured and structured analytics under a single abstraction, by leveraging the immense potential of LLMs to process multi-modal datasets. At the same time, given the limitations of current models, we are building Aryn to be a human-in-the-loop system; as the LLMs improve, the need for human interventions will diminish, but it is unlikely to completely vanish. Our experience across a variety of application domains has validated our overall design as well as Aryn’s individual components. Of course, many challenges still remain. We need more robust data and querying abstractions to support analytics across structured and unstructured repositories. Knowledge Graphs are emerging as a strong contender for this purpose, but it is not yet clear how hallucination-free knowledge graphs can be built in a cost-effective manner with minimum human intervention. We also need to build more scalable systems for processing large multi-modal repositories, for inferring connections between them, and for leveraging existing structured repositories, like data warehouses.